# Toward Explainable Offline RL: Analyzing Representations in Intrinsically Motivated Decision Transformers

### v1

@@ Abstract
Elastic Decision Transformers (EDTs) have proved to be particularly successful in offline reinforcement learning, offering a flexible framework that unifies sequence modeling with decision-making under uncertainty. Recent research has shown that incorporating intrinsic motivation mechanisms into EDTs improves performance across exploration tasks, yet the representational mechanisms underlying these improvements remain unexplored. In this paper, we introduce a systematic post-hoc explainability framework to analyze how intrinsic motivation shapes learned embeddings in EDTs. Through statistical analysis of embedding properties (including covariance structure, vector magnitudes, and orthogonality), we reveal that different intrinsic motivation variants create fundamentally different representational structures. Our analysis demonstrates environment-specific correlation patterns between embedding metrics and performance that explain why intrinsic motivation improves policy learning. These findings show that intrinsic motivation operates beyond simple exploration bonuses, acting as a representational prior that shapes embedding geometry in biologically plausible ways, creating environment-specific organizational structures that facilitate better decision-making.
1 Introduction
In recent years, the field of reinforcement learning (RL) RLReview has undergone an impactful evolution, moving beyond reactive policy optimization to also include models that exhibit larger generalization and adaptability levine2020offline .Reinforcement Learning (RL) trains agents to interact with an environment and learn from rewards. Different types of RL approaches exist, among which Offline Reinforcement Learning (Offline RL), also referred to as batch RL. Such approach aims to learn optimal decision-making policies solely from previously collected data, without further environment interaction during training. This setting is especially relevant in domains where online exploration is costly, risky, or infeasible — such as robotics, healthcare applications and autonomous driving kidambi2020morel ; agarwal2020optimistic ; kumar2020conservative ; prudencio2023survey . Among the several architecture proposed in the field, Elastic Decision Transformers (EDTs), proposed in wu2023elastic , have emerged as a novel and promising architecture, unifying sequence modelling with action-conditioned decision-making. More specifically Elastic Decision Transformers, leveraging the strengths of Transformer architectures (traditionally dominant in natural language processing), can capture efficiently long-range dependencies and enabling flexible policy behaviours under uncertainty. In addition to the more traditional RL paradigms, the paradigms of intrinsic rewards and curiosity driven models. Such models are in fact driven by concepts as for instance intrinsic motivation and curiosity, which can encourage significantly the agents’ exploration and learning experience oudeyer2007intrinsic . Adding intrinsic motivation can in fact enhance potentially the learning experience, adding the capability of a further world exploration driven by the agent in a pro-active way. However, as RL models can be enriched with intrinsic motivation, as shown in esann2025 , it still remains largely unexplored the explainability of the improvement brought by the intrinsic motivation to the RL reinforcement learning models where intrinsic motivation and curiosity driven approaches have been has been added pathak2017curiosity . Understanding how and why an RL intrinsic agent arrives at its decisions represents therefore a central research question, which could lead to even a wider adoption of such models, which could turn out to be critical, especially in safety-critical or exploratory domains. In contrast to traditional RL methods, where hand-crafted features or value estimates often yield some interpretability, EDTs learn implicit state representations within high dimensional embedding spaces. These representations, while powerful, needs deeper interpretation in order to understand and disentangle their internal meaning bengio2013representation . One promising path toward interpretability lies in the study of representation learning bengio2013representation : specifically, how internal embeddings evolve under different learning signals. Intrinsic motivation, a mechanism inspired by cognitive science, has been shown to enhance exploration in sparse-reward environments by encouraging agents to seek novel or surprising states. Recent work has incorporated intrinsic rewards into EDTs, yielding improved performance across a suite of offline RL benchmarks. However, the deeper representational consequences of these signals (the extent to which they influence the structure, semantics, and coherence of learned embeddings) remain largely unexplored, and matter of interest for understanding the different performances obtained by curiosity enhanced RL models. In this paper, we take a step toward opening the black box of EDTs. By applying statistical analysis techniques, we investigate the internal embedding spaces of EDTs trained with and without intrinsic motivation. Our goal is to assess whether intrinsic rewards do more than simply guide the agents’ behavior: does intrinsic motivation also change the geometry of learned latent representations, in a way that supports explainability? Our findings suggest that agents equipped with intrinsic rewards tend to perform better, developing sort of more structured and disentangled representations, pointing toward an emergent semantic alignment between states, goals, and latent dynamics. Through systematic statistical analysis of embedding properties, we demonstrate that intrinsic motivation mechanisms sculpt representation geometry in environment-specific ways that correlate with improved policy learning. These findings can bridge the gap between empirical performance gains and the underlying representational mechanisms, offering new insights into both the effectiveness and interpretability of intrinsic motivation in transformer-based offline RL. More specifically, in our paper we provide, to the best of our knowledge, the first systematic analysis of how intrinsic motivation mechanisms shape learned representations in Elastic Decision Transformers. We propose several novelties and contributions, among which:
@@ 1 Introduction
    Post-Hoc Explainability: We propose a statistical analysis framework (covariance trace, L2 norm, cosine similarity) to examine how intrinsic motivation shapes embedding geometry, revealing differences between EDT models with and without curiosity mechanisms.
@@
    Mechanistic Analysis: We introduce two EDT variants: EDT-SIL, where intrinsic loss acts on embedded states promoting compactness, and EDT-TIL, where it operates on transformer outputs, enhancing orthogonality. These lead to distinct representational structures correlated with environment-specific performance gains.
@@
    Performance-Representation Link: We demonstrate quantitative correlations between embedding properties and task performance across environments, offering mechanistic insights into when and why intrinsic motivation improves offline RL outcomes.
2 Background
Reinforcement learning (RL) has witnessed significant advancements in the latest years, with the integration of several latest models within the learning setting. Among the several approaches proposed, the inclusion of Decision Transformer (DT) models have introduced a significant novelty, by framing RL as a sequence modeling research problem, leveraging the Transformer architecture to predict actions based on past trajectories chen2021decision . However, DTs face challenges in trajectory stitching, that is in the process of combining optimal segments from sub-optimal trajectories, to aim at better policies. In this context, the Elastic Decision Transformer (EDT) model was proposed, enhancing DTs by dynamically adjusting the history length, considered during action inference wu2023elastic . In fact, through a modulation of the input sequence length, based on the quality of past trajectories, EDTs are able to enable effective trajectory stitching, leading to improved performance in the context of offline RL settings. Another important aspect, concerns the possible use of intrinsic signals and motivation within the reinforcement learning paradigms. In fact, in reinforcement learning (RL), agents traditionally learn by maximizing extrinsic rewards—signals provided by the environment to guide behaviour. However, in several real-world environments, such rewards are sparse, delayed, or poorly aligned with long-term success levine2020offline . This feature, has driven towards the development of alternative RL frameworks, based on intrinsic motivation or curiosity driven approaches, a framework that assign to agents some sort of internal drives to explore and learn, even in the absence of external feedbacks pathak2017curiosity . Intrinsic motivation takes inspiration from cognitive psychology, in particular from theories of curiosity and novelty-seeking behaviour, observed in animals and humans oudeyer2007intrinsic ; pathak2017curiosity .Therefore, in the RL setting, intrinsic rewards can be seen as a sort of pseudo-additional reward that encourages exploratory behaviours by rewarding novelty, uncertainty, surprise, or learning progresses. Several previous examples of approaches based on intrinsic motivations can be found. For instance in bellemare2016unifying the authors present the novelty-based methods, based on count-based exploration in large or continuos state spaces. Furthermore in pathak2017curiosity the authors present prediction error-based methods, like Intrinsic Curiosity Modules (ICM), where reward agents when their predictive models fail—encouraging behaviours that reveal new knowledge. In houthooft2016vime instead, the authors present information gain approaches, which measure KL divergences between posterior and prior beliefs to guide the agent towards informative experiences. Recent work has aimed to unify intrinsic and extrinsic drives. In chen2022redeeming the authors propose a constrained optimization framework that balances policy optimization across both reward types, while others use contrastive learning and large-scale embeddings (e.g., CLIP) to define intrinsic rewards based on semantic novelty gupta2022foundation . Despite these advances, aligning intrinsic motivation with downstream task success remains an open challenge. Poorly designed intrinsic rewards can distract agents from task-relevant behaviours or lead to inefficient exploration. Ongoing research continues to investigate better ways to quantify novelty, adjust reward weighting dynamically, and integrate intrinsic objectives with policy learning in a principled manner. In the context of RL and exploration, embeddings can play a pivotal role in RL by providing compact representations of states, actions, and rewards. The use of pre-trained models, such as CLIP, allows agents to leverage semantic knowledge for better generalization. For instance in gupta2022foundation the authors demonstrated that CLIP-based embeddings could serve as intrinsic rewards, guiding agents towards semantically meaningful exploration and outperforming traditional methods in complex environments. In this context we would also like to consider a few works which address the issue of explainable reinforcement learning. The majority of works in this area, focuses in fact on the concept of explainability of the decisions and actions taken by the agent, as reported in the complete survey proposed in milani2024explainable . In our work, instead, the goal is to define explainability in the context of deriving insights on the advantages of the use of the curiosity driven approach using embedded within the EDT framework, disentangling the inherent geometry and characteristics of the learnt embeddings.
2.1 Biological Plausibility and Homeostatic Regulation: towards a biologically inspired reinforcement learning approach
In this section we would like to propose some insights on the biological plausibility of our proposed model. The quest for explainable AI in reinforcement learning (RL) has a tight parallel with biological learning, where intrinsic motivation shapes adaptive behavior through allostatic regulation, achieving stability by adjusting predictions rather than fixing a–priori parameters allostasi1988 . Such predictive adaptation is reflected in how the so called Random Network Distillation (RND) models operate: learning is driven by discrepancies between predicted and actual inputs, echoing brain mechanisms that constantly update internal models based on sensory prediction errors friston2010free ; Clark_2013 . These prediction hierarchies span from primary sensory to higher-order cortical processing friston2009predictive , suggesting intrinsic motivation can be applied across representational levels. Moreover, biological systems maintain representational homeostasis, optimizing information processing through the regulation of capacity and structural organization zenke2017 ; ganguli2012compressed . Intrinsic motivation fosters flexible learning and generalization oudeyer2007intrinsic . In transformer-based models like Elastic Decision Transformers, auxiliary losses based on RND act as allostatic regulators, guiding representational structure without altering offline reward signals. Such losses can help preventing representational collapse, mirroring biological mechanisms that sustain learning adaptability and predictive efficiency.
@@ 3 Methods
Figure 1: Architecture of the Elastic Decision Transformer with intrinsic motivation mechanisms proposed in esann2025 . Figure 1 shows both EDT-SIL and EDT-TIL variants, where the RND module operates on state embeddings or transformer outputs respectively. The dashed lines indicate backpropagation paths for each variant, with $L_{int}$ from the RND module contributing to the total loss $L_{EDT}$ alongside standard prediction losses. The target network of the RND module has frozen weights and is never updated, as proposed in rnd .
In this section, we will provide a comprehensive overview of the Elastic Decision Transformer architecture enhanced with intrinsic motivation mechanisms. While maintaining the same architecture as esann2025 , we provide detailed descriptions of the baseline framework and the two intrinsic motivation variants that form the basis for our embedding analysis presented in this work. More specifically, we extended the work esann2025 by providing a comprehensive evaluation across different datasets. We then further conducted a systematic investigation of RND network layer configurations, and present our post-hoc explainability framework for analyzing how intrinsic motivation shapes learned representations, towards obtaining better final performances.
3.1 Intrinsic Auxiliary Loss
Our approach builds upon the Elastic Decision Transformers models proposed in wu2023elastic , which excel in offline RL by processing trajectories as sequences of (state, action, reward) tuples and predicting actions conditioned on states and desired returns. The baseline EDT leverages Transformer attention mechanisms to capture long-range dependencies in sequential decision-making. Following esann2025 , we analyze the intrinsically-motivated EDT variants that incorporate an RND module as an auxiliary loss function, operating independently of fixed reward signals in offline datasets. As shown in Figure 1, these variants include:
  • EDT-SIL (State Input Loss): In this variant, the RND module computes intrinsic rewards directly from embedded state representations. The predictor network attempts to match the output of a frozen target network when fed the embedded states. This configuration allows the intrinsic signal to influence the state embedding layer through backpropagation, potentially encouraging the model to learn more diverse and structured state representations.
  • EDT-TIL (Transformer Input Loss): Here, the RND module operates on the transformer’s output representations, aligning intrinsic reward computation with the model’s sequential processing capabilities. This approach enables the intrinsic signal to shape both the embedding and transformer layers, potentially creating more coherent sequential representations.
The intrinsic loss $L_{\text{int}}$ is computed as the mean squared error between the predictor and target network outputs as described in Equation 1:
$L_{\text{int}}=|f_{\text{pred}}(x;\theta_{\text{pred}})-f_{\text{target}}(x;\theta_{\text{target}})|_{2}^{2}$ (1)
@@ 3.1 Intrinsic Auxiliary Loss
where $x$ represents either embedded states (SIL) or transformer outputs (TIL). The total loss is computed summing up the standard EDT objective with this intrinsic component as described in Equation 2:
$L_{\text{overalll}}=L_{\text{EDT}}+L_{\text{int}}$ (2)
For the complete formulation of $L_{\text{EDT}}$, including its constituent components (return prediction, observation prediction, action prediction, and expectile regression losses), we refer the reader to wu2023elastic . This formulation enables the intrinsic motivation signal to enhance representation learning without disrupting the primary task objective or altering the fixed reward structure of offline datasets.
3.2 Random Network Distillantion (RND) Architecture Analysis
As described in Figure 1 we introduce the RND block. In order to understand the impact of RND network capacity on intrinsic motivation effectiveness considering network depth, we conducted a systematic investigation of the predictor network depth. The motivation for this analysis stems from the hypothesis that different network capacities may capture different levels of representational complexity, potentially affecting both the quality of intrinsic rewards and the resulting policy performance. We evaluated in total 3 RND predictor configurations: a 1-layer RND (considering it as a minimal architecture to establish a baseline for intrinsic reward generation), a 3-layers RND as the default configuration from prior work esann2025 , and a 10-layers RND, considering it as a high-capacity variant to test whether increased expressiveness improves intrinsic motivation. This analysis was conducted on our best performing dataset (the Medium Datasets as described in Section 4.1), in order to optimize the RND predictor network architecture.
3.3 Post-Hoc Explainability Analysis
The primary methodological contribution of this work lies in developing a framework for post-hoc explainability, in particular evaluating and studying how intrinsic motivation mechanisms shape learned representations. This analysis aims to understand why intrinsic motivation improves performance, by examining the geometric and statistical properties of embedding spaces.
3.3.1 Embedding Characterization Framework
  • Covariance Trace: This metric measures the total variance distributed across embedding dimensions as per Equation 3:
    $cov\_trace=\mathrm{Tr}(\mathrm{Cov}(E))$ (3)
  • L2 Norm: The mean magnitude of embedding vectors quantifies representational compactness, as per Equation 4:
    $l2\_norm=\frac{1}{N}\sum_{i=1}^{N}|e_{i}|_{2}$ (4)
  • Cosine Similarity: The average pairwise cosine similarity within each tensor assesses representational orthogonality as per Equation 5:
3.3.2 Performance-Representation Correlation Analysis
To establish quantitative relationships between representational properties and task performance, we first used correlation analysis between embedding metrics and normalized performance scores (HNS defined in Equation 6). This approach allowed us to identify which representational characteristics are most predictive of policy performance in each environment. Our methodology included:
  1. 1.
    Computing embedding metrics for each trained model across multiple seeds
  2. 2.
    Computing the Pearson correlations between metrics and performance scores
  3. 3.
    Identifying the most predictive metric for each environment-model combination
  4. 4.
    Making an analysis on how different intrinsic motivation mechanisms (EDT-SIL vs. EDT-TIL) can create distinct representational patterns
Such post-hoc analysis allowed us to provides insights into the mechanistic basis for performance improvements observed with intrinsic motivation, moving beyond empirical results to understand the underlying representational changes that enable better policy learning.
@@ 4 Experiments and Results
In this section, we will provide a comprehensive evaluation of intrinsic motivation mechanisms in Elastic Decision Transformers, focusing on both performance improvements and the underlying representational changes. We use the standard EDT architecture described in wu2023elastic as our baseline, comparing it against the intrinsic motivation variants (named EDT-SIL and EDT-TIL and described in Section 3.1). Our experimental design systematically investigates how these variants shape learned embeddings and their correlation with task performance across multiple locomotion environments as described in Section 4.1.
4.1 Experimental Setup and Datasets
Figure 2: The four locomotion environments from D4RL fu2020d4rl benchmark: Hopper (top-left), Walker2d (top-right), HalfCheetah (bottom-left), and Ant (bottom-right). These tasks represent different locomotion challenges from bipedal to quadrupedal movement.
We conducted experiments on four continuous control tasks from the D4RL benchmark suite: Ant, HalfCheetah, Hopper, and Walker2d (Figure 2). D4RL (Datasets for Deep Data-driven Reinforcement Learning and presented in fu2020d4rl ). These dataset provide a standardized collection of offline RL datasets, that address critical challenges in offline learning by offering diverse datasets with varying quality, coverage, and collection methodologies. The different locomotion tasks represent different movement challenges, which are particularly relevant for evaluating biological plausibility. The tasks range from quadrupedal movement (Ant) to bipedal locomotion with varying degrees of stability (Hopper, Walker2d) and high-speed running (HalfCheetah). Each environment provides distinct sensorimotor dynamics that mirror the variety of locomotion patterns found in biological systems, making them ideal testbeds for intrinsic motivation mechanisms inspired by natural learning processes.
For each environment, we evaluated our models on both medium and medium-replay datasets. The difference between the two is the following: medium datasets ($\sim$1M transitions) provide cleaner trajectories from behavior policy training, capturing progression from suboptimal to near-optimal performance. Medium-replay datasets ($\sim$2M transitions) add noisy replay buffer data including early exploration, creating more challenging conditions that better reflect real-world learning scenarios. Such datasets are ideal for investigating biological plausibility as they mirror natural motor skill development through trial and error, making them suitable for understanding how intrinsic motivation shapes representations in biologically plausible ways.
All experiments were conducted using five random seeds to ensure statistical robustness. Performance was evaluated using Human-Normalized Scores (HNS), calculated as:
@@ 4.1 Experimental Setup and Datasets
This metric provides consistent scaling across environments and follows the same evaluation protocol used in wu2023elastic . Each trained model underwent evaluation over three rounds of 100 episodes each, with results averaged across both seeds and evaluation rounds.
For the embedding analysis, we collected state embeddings obtained during the model evaluation steps, by executing the best performing model for each environment-dataset combination. During the evaluation phase, models ran for a single episode with a maximum of 1000 steps. At each step, we extracted and saved the state embeddings, i.e. the output of the embedder layer as shown in Figure 1. These embeddings represent the agent’s internal representation of environmental states and served as the input to the RND module in the EDT-SIL variant. The choice to analyze state embeddings was motivated by biological plausibility: these representations correspond to the primary sensory encoding stage in biological neural systems, where environmental observations are first transformed into internal neural patterns ganguli2012compressed . Similarly to biological systems must efficiently encode sensory information before higher-order processing, the embedder layer can, in fact, create the foundational representations upon which all subsequent decision-making depends. From these collected embeddings, we computed 3 key geometric metrics (covariance trace, L2 norm, and cosine similarity) following the analysis framework detailed in our Section 3. The embedding collection and metric calculation process was repeated three times to ensure robustness, with all results reported in Table 1 representing averaged values across these repetitions.
4.2 RND Layer Configuration Analysis
To optimize the intrinsic motivation mechanism, we systematically investigated the effect of RND predictor network depth on both performance and representation quality. We evaluated three configurations: 1-layer, 3-layer, and 10-layer RND networks. This analysis was conducted exclusively on medium datasets, as these demonstrated the most promising initial results. The 3-layer configuration emerged as optimal across both EDT-SIL and EDT-TIL variants. Figure 3 demonstrates that 3-layer variants (highlighted with red borders) consistently achieve the highest cumulative scores across all environments. This finding aligns with the biological principle of representational balance: too few layers (1-layer) may lack sufficient capacity to capture complex predictive relationships, while too many layers (10-layer) may lead to overfitting or representational instability.

### v2

# How Intrinsic Motivation Shapes Learned Representations in Decision Transformers: A Cognitive Interpretability Analysis

@@ Abstract
Elastic Decision Transformers (EDTs) with intrinsic motivation have demonstrated improved performance in offline reinforcement learning, yet the cognitive mechanisms underlying these improvements remain unexplored. We introduce a systematic post-hoc explainability framework to analyze how intrinsic motivation shapes learned embeddings in EDTs through statistical analysis of embedding properties (covariance structure, vector magnitudes, and orthogonality). We reveal that different intrinsic motivation variants create fundamentally different representational structures: one variant operating on state embeddings promotes compact representations, while another operating on transformer outputs enhances representational orthogonality. Our analysis demonstrates strong environment-specific correlation patterns between embedding metrics and performance across locomotion tasks. These findings show that intrinsic motivation operates as a representational prior that shapes embedding geometry in cognitively plausible ways, creating environment-specific organizational structures that facilitate better decision-making beyond simple exploration enhancement.
1 Introduction and Background
Reinforcement learning has evolved beyond reactive policy optimization to include models with greater generalization and adaptability, particularly in offline reinforcement learning where agents learn optimal policies solely from previously collected data without environment interaction during training levine2020offline; kidambi2020morel; kumar2020conservative. Elastic Decision Transformers (EDTs) wu2023elastic have emerged as a promising architecture that unifies sequence modeling with decision-making by leveraging Transformer transformer architectures to capture long-range dependencies and enable flexible policy behaviors under uncertainty. EDTs enhance standard Decision Transformers chen2021decision through dynamic history length adjustment, enabling effective trajectory stitching for improved offline RL performance.
Intrinsic motivation mechanisms, inspired by cognitive science theories of curiosity and novelty-seeking behavior oudeyer2007intrinsic; pathak2017curiosity, have been incorporated into RL to encourage exploration when extrinsic rewards are sparse or poorly aligned with long-term success. Recent work has shown that integrating intrinsic rewards into EDTs yields improved performance across offline RL benchmarks esann2025. However, while the performance benefits are empirically established, the representational mechanisms underlying these improvements remain largely unexplored. Understanding how intrinsic motivation shapes the internal embedding spaces of EDTs is crucial for interpretable reinforcement learning, as these models learn implicit state representations in high-dimensional spaces that lack the interpretability of traditional hand-crafted features bengio2013representation.
This paper addresses the gap between empirical performance gains and mechanistic understanding by investigating how intrinsic motivation shapes learned representations in EDTs. We introduce a systematic post-hoc explainability framework using statistical analysis of embedding properties (covariance trace, L2 norm, cosine similarity) to examine representational geometry. Our analysis reveals that intrinsic motivation operates beyond simple exploration bonuses, acting as a representational prior that creates environment-specific organizational structures. We analyze the two EDT variants presented in esann2025: EDT-SIL, where intrinsic loss operates on embedded states promoting compactness, and EDT-TIL, where it operates on transformer outputs enhancing orthogonality. Our contributions include:
@@ 1 Introduction and Background
    Post-hoc explainability framework for analyzing embedding geometry changes.
@@
    Mechanistic analysis revealing distinct representational structures of EDT variants.
@@
    Demonstration of quantitative correlations between embedding properties and task performance across environments.
@@ 2 Methods
We analyze Elastic Decision Transformers enhanced with intrinsic motivation mechanisms esann2025, building upon the EDT architecture wu2023elastic that processes trajectories as sequences of (state, action, reward) tuples. Our analysis focuses on two intrinsically-motivated EDT variants that incorporate Random Network Distillation (RND) rnd modules as auxiliary loss functions.
2.1 Intrinsic Motivation Variants
We examine two EDT variants that differ in where the intrinsic signal operates: EDT-SIL (State Input Loss) computes intrinsic loss directly from embedded state representations, allowing the intrinsic signal to influence the state embedding layer and potentially encourage more structured representations. EDT-TIL (Transformer Input Loss) operates on transformer output representations, enabling the intrinsic signal to shape both embedding and transformer layers for more coherent sequential representations.
@@ 2.1 Intrinsic Motivation Variants
The intrinsic loss is computed as $L_{\text{int}}=|f_{\text{pred}}(x;\theta_{\text{pred}})-f_{\text{target}}(x;\theta_{\text{target}})|_{2}^{2}$ where $x$ represents either embedded states (SIL) or transformer outputs (TIL). The total loss combines the standard EDT objective with this intrinsic component: $L_{\text{overall}}=L_{\text{EDT}}+L_{\text{int}}$. This formulation enables intrinsic motivation to enhance representation learning without disrupting the primary task objective.
2.2 Post-Hoc Explainability Framework
Our primary contribution is a framework for analyzing how intrinsic motivation shapes learned representations by examining geometric and statistical properties of embedding spaces. We focused on three key metrics that capture different aspects of representational structure: covariance trace measuring total variance distribution across embedding dimensions, L2 norm quantifying representational compactness, and cosine similarity assessing representational orthogonality.
To establish quantitative relationships between representational properties and task performance, we computed Pearson correlations between embedding metrics and normalized performance scores across multiple seeds, identifying the most predictive metric for each environment-model combination. This analysis reveals how different intrinsic motivation mechanisms create distinct representational patterns and provides mechanistic insights into why intrinsic motivation improves policy learning.
@@ 3 Experiments and Results
We evaluate intrinsic motivation mechanisms in Elastic Decision Transformers, focusing on performance improvements and underlying representational changes. Using the standard EDT architecture wu2023elastic as baseline, we compare against EDT-SIL and EDT-TIL variants across four continuous control tasks from the D4RL benchmark fu2020d4rl: Ant, HalfCheetah, Hopper, and Walker2d. These locomotion tasks represent different movement challenges from quadrupedal (Ant) to bipedal locomotion (Hopper, Walker2d) and high-speed running (HalfCheetah), providing diverse sensorimotor dynamics for evaluating intrinsic motivation mechanisms.
We evaluated models on both medium datasets ($\sim$1M transitions with cleaner trajectories) and medium-replay datasets ($\sim$2M transitions with noisy replay buffer data), using five random seeds for statistical robustness. Performance was evaluated using Human-Normalized Scores (HNS), providing consistent scaling across environments:
@@
For embedding analysis, we collected state embeddings during evaluation by executing the best performing model for each environment-dataset combination, extracting embeddings at each step over single episodes with maximum 1000 steps. From these embeddings, we computed three key geometric metrics (covariance trace, L2 norm, and cosine similarity) following our analysis framework, with results averaged across three repetitions for robustness.